# Tunable transport in bi-disperse porous materials with vascular structure

### v1

@@ IV.2 Transpiration
Figure 5 presents our measurements of transpiration flow rates for each sample geometry as a function of water potential (Equation 21) of the vapor to which the sample was exposed. After each step in vapor pressure, we observed clear steady-state flows (as measured by the emptying speed, $U=\mathrm{d}X/\mathrm{d}t$ of the serpentine channel) through the structure, after short transients of a few minutes. Lowering $p$ below $p_{\mathrm{sat}}$ increased the magnitude of the water potential $\Psi(p)$, controlling the negative pressure driving the flow to the open edge. As seen in Figure 5c, we observed for each sample a linear relationship between the flow rate and $\Psi(p)$ (i.e. a logarithmic dependence of $Q$ on $p$), as predicted by Equation (20)-(21). We extracted permeability ratios, $\kappa/\kappa_{0}$ (see section III.3) using linear fittings of the experimental data and Equations (1) and (20). See below for a general discussion of these results (section IV.4 and Figure 7b).
@@ IV.4 Global results
Figure 7 summarizes the effective permeability measurements obtained from numerical simulations, analytical estimates, and from all the experimental approaches developed above. Figure 7a compares the approximate, analytical estimate (Equation 17) to numerical simulations results, and is a graphical representation of the rightmost two columns in Table 1. While we use finite difference simulations as a reference, it is sometimes more convenient to use simpler approaches for design rules-of-thumb. As can be seen in Figure 7a (and more generally in Figure 4), Equations (17-17) provide reasonable estimates of what can be expected for a particular geometry, with typically less than $20\%$ error in $\kappa/\kappa_{0}$ when the surface coverage of the microchannels exceeds $0.3$.

### v2

@@ IV.2 Transpiration
Figure 6 presents our measurements of transpiration flow rates for each sample geometry as a function of water potential (Equation 21) of the vapor to which the sample was exposed. After each step in vapor pressure, we observed clear steady-state flows (as measured by the emptying speed, $U=\mathrm{d}X/\mathrm{d}t$ of the serpentine channel) through the structure, after short transients of a few minutes. Lowering $p$ below $p_{\mathrm{sat}}$ increased the magnitude of the water potential $\Psi(p)$, controlling the negative pressure driving the flow to the open edge. As seen in Figure 6c, we observed for each sample a linear relationship between the flow rate and $\Psi(p)$ (i.e. a logarithmic dependence of $Q$ on $p$), as predicted by Equation (20)-(21). We extracted permeability ratios, $\kappa/\kappa_{0}$ (see section III.4) using linear fittings of the experimental data and Equations (1) and (20). See below for a general discussion of these results (section IV.4 and Figure 8b).
@@ IV.4 Global results
Figure 8 summarizes the effective permeability measurements obtained from numerical simulations, analytical estimates, and from all the experimental approaches developed above. Figure 8a compares the approximate, analytical estimate (Equation 17) to numerical simulations results, and is a graphical representation of the rightmost two columns in Table 1. While we use finite difference simulations as a reference, it is sometimes more convenient to use simpler approaches for design rules-of-thumb. As can be seen in Figure 8a (and more generally in Figure 5), Equations (17-18) provide reasonable estimates of what can be expected for a particular geometry, with typically less than $20\%$ error in $\kappa/\kappa_{0}$ when the surface coverage of the microchannels exceeds $0.3$.